# LITMUS: Bayesian Lag Recovery in Reverberation Mapping with Fast Differentiable Models

### v1

@@ 6 Discussion & Future Work
LITMUS’s improved handling of the aliasing problems of multi-year RM surveys means that, by being applied to existing data, it stands to improve the statistical power of our RM results without the need for new observations. OzDES-McDougall_2025 found that the collective constraining power of the entire literature of RM measurements is sufficient to constrain the high redshift $R-L$ relationships such that the statistical uncertainty is subdominant compared to the inherent population noise. However, it is not confirmed whether a single $R-L$ relationship is sufficient to explain all AGN behaviour over all time. Current constraints are not enough to test the possibility of time-varying $R-L$ relationships, and the existing sample is restricted to only a narrow window of redshift-luminosity space. Through LITMUS’s hypothesis testing we can not only successfully recover more lags, but also expand their breadth in parameter space, and so will be equipped to begin using RM as a probe of more expressive physical models.
@@
LITMUS is a general purpose single-lag RM tool, and so can be applied beyond this BLR RM domain. The same GP statistical model can be applied to quasar time delay without much alteration (e.g. see the GP based approaches in Ding_2021_TDLMC2)), as well as any other single-lag estimation applications. Because LITMUS’s statistical models can be extended or swapped out, it can also aimed at answering some of the stickier questions of BLR RM, such as means of accounting for the BLR holiday (Dehghanian_2019) or descriptions of outlier measurements.
@@ Acknowledgements
The author would like to thank Louis Desdoigts for his experience and advice in writing LITMUS, and Rahma Alfarsy of the DESI collaboration for aiding in testing its initial stages. All software was made using python (VanRossum_2009_python) and with the aid of numpy (harris_2020_numpy). Plots and figures were generated with the aid of matplotlib (Hunter_2007_matplotlib) and chainconsumer (Hinton_2016_chainconsumer).
@@
We acknowledge the traditional custodians of the Macquarie University land, the Wallumattagal clan of the Dharug nation, whose cultures and customs have nurtured and continue to nurture this land since the Dreamtime. We pay our respects to Elders past, present and future.
@@
Appendix A The Grid Smoothing Algorithm
The Laplace Quadrature and SVI Quadrature algorithms divide the lag domain into an uneven grid of Gaussian slices. If these are spaced too sparsely we can miss important information about the posterior peaks, particular in the narrowly constrained modes of high SNR cases, and if they are packed too densely we can lose information about the long tails of the distribution which can impact the accuracy of the evidence estimate (see Figure A). In the the idealised case, the two extremes are to have the points evenly spaced across the entire domain ($\delta\Delta t$ is constant) and having them distributed with spacing inversely proportional to the posterior density ($\delta\Delta t\propto P(\Delta t)$).
In LITMUS we describe a sliding scale between these with a “grid bunching” parameter, $\alpha$, where $\alpha=0$ is an evenly spaced grid and $\alpha=1$ is bunched proportional to the posterior. Values in between describe spacings whose cumulative distribution function (CDF) is a weighted average of these two extremes. First, we find some estimate of the best-fit parameters for the entire model, $\hat{\theta}$, and from this point forward fix all non-lag parameters at these values, $\hat{\phi}$, to produce a conditional distribution $\propto P(\Delta t|\hat{\phi})$. Estimating the entire marginal distribution $P(\Delta t)$ would be equivalent to solving the entire distribution, and so we form the grid based only on estimates of this much simpler conditional distribution.
For brevity, define $x=\frac{\Delta t-\Delta t_{\mathrm{Min}}}{\Delta t_{\mathrm{Max}}-\Delta t_{\mathrm{Min}}}$ and function $y(x)=\mathcal{P}(\Delta t(x)|\hat{\phi})$. The grid smoothing algorithm is as follows:
@@ Appendix A The Grid Smoothing Algorithm
[…]
@@
Once converged (in Figure A this is after $5$ iterations), we construct a grid of $I$ ordered $x$ values from the final $Y(x)$ and convert to the desired set of lags ${\Delta t^{i}}$ with $\Delta t^{i}=\Delta t_{\mathrm{Min}}+x_{i}^{J-1}(\Delta t_{\mathrm{Max}}-\Delta t_{\mathrm{Min}})$.
@@
Appendix B Optimisation & Preconditioning
@@
$\mu_{\mathrm{cont/resp}}$ $=\left(\sum_{i}E_{i}^{-2}\right)^{-1/2}\sum_{i}{E_{i}^{-2}y_{i}},$ (25)
$\sigma_{\mathrm{cont/resp}}$ $=\left(\sum_{i}E_{i}^{-2}\right)^{-1/2}\sum_{i}{E_{i}^{-2}(y_{i}-\mu)^{2}}.$
To find a good starting value for $\tau$, we use the ICCF method outlined in Section 3.2 to estimate the auto-covariance function of the continuum. This estimate becomes increasingly less accurate at high autocorrelation times $\delta t$, but drawing some window $\lvert\delta t\rvert<a$ it should follow close to the exponential form in Equation 9. Knowing this, we can say that $\ln\left|\mathrm{ACF}(\delta t)\right|\approx-\delta t/\tau+2\ln\left|\sigma\right|$. Drawing some window $\lvert\delta t\rvert<a$ and inverting the $\delta t>0$ side, this logarithm becomes roughly linear with slope $\frac{\partial y}{\partial\delta t}$. Taking a standard linear regression of this function, we can then estimate $\tau$. Precision is not too important here, as the model parameter is $\ln\left|\tau\right|$ in GP fitting and so is insensitive to small deviations.
@@
    Exhibit a severe drop in peak the posterior density compared to the previous slice, which would suggest the slice is in one of the furrows shown in Figure 2.
@@
    An approximate form of (ii) where we take only the diagonal elements of $H$, i.e. no skewing / rotating of the axes.
For stability purposes, any negative or zero eigenvalues are fixed to unity in approaches (ii) and (iii).

### v2

@@ 4.4 Hypothesis Testing
5 Validation of Results
The example in Section 4.4 demonstrates LITMUS’s success on a small number of mocks, but it is necessary to also show that it performs reliably at scale. In this section, we test and compare LITMUS’s performance against existing methods on three sets of mocks:
  1. 1.
    A set of $440$ mock AGN with lags distributed uniformly in $\Delta t\in[0,1000]\;\mathrm{d}$ and timescales drawn log-uniformly in $100-1000\;\mathrm{d}$.
  2. 2.
    A similar sample of $490$ mocks with lags drawn from a log-normal distribution, $\Delta t\sim\mathcal{N}(2,0.4)$, roughly emulating the spread of the OzDES Mgii sample.
  3. 3.
    The same set of $490$ mocks in set 2 but with the response curves randomly regenerated to create lag-free mocks, similar to the decoupled responses in Fig 6.
Mock set 1 allows us to interrogate LITMUS’s performance in lag recovery and its resistance to aliasing, particularly in contrast to JAVELIN’s AEIS method, while mock sets 2 and 3 allow us to examine how the Bayes factor allows us to screen for false positive lags. The decoupled mocks are of physical interest as they correspond to the case of the lag being longer upper limit detectable in a particular survey, as can happen for longer observer-frame Civ lags (OzDES-Penton_2025). Mocks are all generated with continuum and response amplitudes of unity and zero mean, with weekly cadence and $\approx 1\%$ uncertainty for the continuum observations and monthly cadence with $\approx 10\%$ uncertainty for the response light curves.
For each mock in each set, we recover the lag with LITMUS’s Laplace Quadrature, SVI Quadrature and Nested Sampling fitting methods, as well as its in-built ICCF and JAVELIN-like AEIS fitting methods. [^3] The AEIS method does not map exactly to a true JAVELIN fit as no smoothing is applied from a transfer function and fitting is performed in the unconstrained domain (See Appendix A). The fitting parameters for each method are listed in Appendix C. For a fair comparison we only include mocks for which all five fitting methods are have fully converged. For the Nested Sampling method we use $10^{4}$ live points, enough to ensure good performance, so that this well established algorithm can act as a benchmark to test the other four fitting methods against. In this benchmarking, use the rough metric of a lag being correct or incorrect if it agrees / differs from the true lag by a threshold of $30\;\mathrm{d}$.
As can be seen in Figure 7, LITMUS’s aliasing friendly fitting methods produce posteriors that map closely to the true mock lags save a few outliers that can be mostly screened with significance tests (see Table 2). We can also see that the Laplace Quadrature’s assumption of Gaussianity does not significantly disrupt this result as compared to Nested Sampling. Conversely, the results from the JAVELIN-like AIES are heavily obscured by the impacts of aliasing, with more than half the posterior samples sitting in the seasonal striations of the aliasing bands. Even before screening spurious lags with Bayes factor tests, LITMUS’s various fitters are significantly better than the AEIS at identifying correct lags.
Figure 7: A comparison of the posterior distributions for the lag error, i.e. the difference between true and recovered lag, comparing some of LITMUS’s aliasing-friendly methods, namely Nested Sampling (left panel) and the Laplace Quadrature (middle panel), to the JAVELIN-like AEIS (right panel). These plots are for mock sample 1 which has $440$ mocks with true lags distributed uniformly over the range $\Delta t\in[0,1000]\;\mathrm{d}$. The aliasing fraction is the fraction of samples / posterior density that sits more than $30\;\mathrm{d}$ from the true value for these mocks. The Laplace Quadrature and Nested Sampling results adhere extremely closely to the true lags save for a single errant false positive, and the similarity between the two validates the Laplace Quadrature’s recovery of the true posterior shape. Conversely, more than half the AEIS samples are incorrect, and the posterior median (black dots) is often far from the true value.
In Figure 8, bottom panel we can see how the Bayes factor acts as a measure of the reliability of lag recoveries. As the Bayes factor between the models of lag-bearing and decoupled response becomes smaller, the error in the recovered lag (measured from the posterior median) rapidly increases. Figure 8 also shows in its top panel how the Bayes factor helps to separate signals that do and do not encode lags within our prior range, with the decoupled population (mock set 3) sitting almost entirely at Bayes factors of $Z_{2}/Z_{1}<1$. This shows how the Bayes factor offers a direct and easily implementable means screening low confidence lag recoveries, and how this quality cuts can be tuned to a nominal level of reliability.
Figure 8: A demonstration of how the Bayes factor acts as a measure of lag reliability. The top panel shows histograms of the Bayes factor evidence ratios for the decoupled mocks with no lag (grey), mocks with a lag that was successfully recovered (navy) and mocks that had an underlying lag but for which the posterior median of the recovery was more than $30\;\mathrm{d}$ from the ground truth. The bottom panel shows how the error in the lag (here, the deviation between ground truth and posterior median) decreases for strong Bayes factors. As the evidence ratio for the lag and decoupled models lowers, the error in the median recovered lag rapidly increases, and above some reasonable threshold (e.g. $Z_{2}/Z_{1}>10^{2}$, the results become significantly more reliable. The solid, dashed and dotted lines represent evidence ratio thresholds of 1:11{:}1, 1:101{:}10 and 1:1001{:}100 in favour of a lag. The correct and incorrect lags are for the $490$ realistic mocks in mock set 2, while the mocks with no lag are from the decoupled mock set 3.
We compare all of LITMUS’s fitting methods, in concert with a quality cut of enforcing $Z_{2}/Z_{1}\geq 10$, to the existing methods of lag recovery (JAVELIN-like AEIS and PyCCF-like ICCF) in Table 2. To roughly emulate the quality cuts of the Mgii OzDES sample, we employ the first $2$ of OzDES-Yu_2023’s quality cuts on the AEIS results, namely that the width of the AEIS posterior, as measured between the 16th and 84th percentiles, is less than $110\;\mathrm{d}$, and that the AEIS and ICCF methods agree to within $2\sigma$, where $\sigma$ is the standard deviation of the AEIS posterior. For the ICCF results we retain only results that agree with the AEIS to within $100\;\mathrm{d}$, and for which the standard deviation of the bootstrapped lag recoveries are also less than $100\;\mathrm{d}$.
Prior to applying cuts, LITMUS’s alias-friendly methods identify the correct lag in both the uniform and log-normal samples with a False Positive Rate (FPR) of $5\%$, while the ICCF and AEIS methods recover the incorrect lag at an FPR of $\approx 40\%$ or higher in these mock sets. After discarding sources with a Bayes factor $Z_{2}/Z_{1}<10$, the LITMUS results improve even further, and importantly retain the vast majority of true recoveries. By contrast the AEIS and ICCF methods, while still seeing a decrease in FPR, do so at the cost of discarding much of the sample. Similarly in the entirely decoupled mock set 3, which contains only false positives, the Bayes factor test removes $\approx 98-99\%$ of these sources while the ICCF and AEIS cuts retain almost ten times more. In short, LITMUS’s methods identify the correct lag more often and performs much better at discarding incorrect lags.
[…]
FPR Before Cuts FPR After Cuts Retained Sources
Mock Set 1 - Uniform Lags
Laplace Quadrature $4.32\%$ $0.57\%$ $80.23\%$
SVI Quadrature $4.09\%$ $0.59\%$ $77.27\%$
Nest. Samp. $3.18\%$ $0.00\%$ $52.05\%$
JAVELIN-Like AIES $41.59\%$ $12.90\%$ $14.09\%$
ICCF $43.41\%$ $30.34\%$ $20.23\%$
Mock Set 2 - Log-Normal Lags
Laplace Quadrature $4.29\%$ $1.30\%$ $94.08\%$
SVI Quadrature $2.45\%$ $1.07\%$ $95.10\%$
Nest. Samp. $2.24\%$ $0.00\%$ $79.39\%$
JAVELIN-Like AIES $66.73\%$ $61.54\%$ $79.39\%$
ICCF $39.39\%$ $18.75\%$ $9.80\%$
Mock Set 3 - Decoupled Response
Laplace Quadrature - - $0.98\%$
SVI Quadrature - - $0.78\%$
Nest. Samp. - - $1.37\%$
JAVELIN-Like AIES - - $13.28\%$
ICCF - - $8.01\%$
Table 2: Summary of the performance of LITMUS’s three fitting methods, the Laplace Quadrature, SVI Quadrature and Nested Sampling approaches, as compared the AEIS fitting method used by JAVELIN and the ICCF method use by PyCCF, as tested on the three sets of mock light-curves (uniform in lag, log-normal in lag to emulate the OzDES $\mathrm{MgII}$ sample, and mocks with the response light curve decoupled such that no true lag is present). Listed are the fraction of false positive lags before and after quality cuts (a lag here being considered incorrect if it differs by more than 30 days from the ground truth), as well as the total number of retained sources after cuts. In general, the LITMUS fitting methods perform significantly better at identifying a true lag where it exists, with a pre-cut $\mathrm{FPR}<5\%$ in all cases, reducing by a factor of a few when removing sources with a lag recovery evidence ratio $Z_{2}/Z_{1}<10$. Overall, LITMUS yields significantly more and and significantly more accurate lags, while also retaining $10-20\times$ fewer spurious lags from the decoupled sample.
@@ 6 Discussion & Future Work
LITMUS’s improved handling of the aliasing problems of multi-year RM surveys means that, by being applied to existing data, it stands to improve the statistical power of our RM results without the need for new observations. OzDES-McDougall_2025 found that the collective constraining power of the entire literature of RM measurements is sufficient to constrain the high redshift RLR-L relationships such that the statistical uncertainty is subdominant compared to the inherent population noise. However, it is not confirmed whether a single RLR-L relationship is sufficient to explain all AGN behaviour over all time. Current constraints are not enough to test the possibility of time-varying RLR-L relationships, and the existing sample is restricted to only a narrow window of redshift-luminosity space. Through LITMUS’s hypothesis testing we can not only successfully recover more lags, but also expand their breadth in parameter space, and so will be equipped to begin using RM as a probe of more expressive physical models.
@@
LITMUS is a general purpose single-lag RM tool, and so can be applied beyond this BLR RM domain. The same GP statistical model can be applied to lensed quasar time delay without much alteration (e.g. see the GP based approaches in Ding_2021_TDLMC2), as well as any other single-lag estimation applications. Because LITMUS’s statistical models can be extended or swapped out, it can also be aimed at answering some of the stickier questions of BLR RM, such as means of accounting for the BLR holiday (Dehghanian_2019) or descriptions of outlier measurements.
@@ Acknowledgements
The authors would like to thank Louis Desdoigts for his experience and advice in writing LITMUS, and Rahma Alfarsy of the DESI collaboration for aiding in testing its initial stages. We thank also Madeline L. Cross-Parkin for her assistance in formatting the plots in this paper and LITMUS’s visual style. All software was made using python (VanRossum_2009_python) and with the aid of numpy (harris_2020_numpy). Plots and figures were generated with the aid of matplotlib (Hunter_2007_matplotlib) and chainconsumer (Hinton_2016_chainconsumer).
@@
We acknowledge the traditional custodians of the Macquarie University land, the Wallumattagal clan of the Dharug nation, whose cultures and customs have nurtured and continue to nurture this land since the Dreamtime. We pay our respects to Elders past and present.
Appendix A The Constrained & Unconstrained Domain
Many Bayesian fitting algorithms encounter difficulties with hard boundaries on model priors, and perform best when parameters are “unconstrained”, i.e. with a domain of $\theta\in\mathbb{R},\;-\infty<\theta<\infty$. Such hard boundaries have infinite gradients in log-probability and other pathologies that can lead to artefacts or non-convergence with some fitting methods. This is a particular concern in RM, where convention is to use bounded uniform priors on all parameters. numpyro obviates these difficulties with a statistical slight of hand. Prior to fitting, all parameters with constrained priors are shifted with a change of variables into a new coordinate space in which they are unconstrained, e.g. for a Uniform distribution on $\theta\sim U_{a,b}(\theta)$ there’s a mapping $f:(a,b)\rightarrow\mathbb{R},\;\theta^{\prime}=f(\theta)$.
$P(\theta)=P(\theta)\times{J(\theta)},$ (20)
where $J(\theta)$ is the Jacobian of the transformation $\theta\rightarrow\theta^{\prime}$. LITMUS’s Laplace Quadrature and SVI Quadrature fitting methods (see Section 4.3) are all performed in this coordinate system by default, i.e. they assume that the posterior is approximately Gaussian in this unconstrained domain. Strictly speaking all densities should be written in terms of $\phi^{\prime}$ and $\theta^{\prime}$ in the description of these algorithms, but we exclude this for simplicity of reading.
Appendix B The Grid Smoothing Algorithm
The Laplace Quadrature and SVI Quadrature algorithms divide the lag domain into an uneven grid of Gaussian slices. If these are spaced too sparsely we can miss important information about the posterior peaks, particular in the narrowly constrained modes of high SNR cases, and if they are packed too densely we can lose information about the long tails of the distribution, which can impact the accuracy of the evidence estimate (see Figure B-9). In the the idealised case, the two extremes are to have the points evenly spaced across the entire domain ($\delta\Delta t$ is constant) and having them distributed with spacing inversely proportional to the posterior density ($\delta\Delta t\propto P(\Delta t)$).
@@ Appendix B The Grid Smoothing Algorithm
For brevity, define
$x=\frac{\Delta t-\Delta t_{\mathrm{Min}}}{\Delta t_{\mathrm{Max}}-\Delta t_{\mathrm{Min}}}$ (21)
[…]
@@
    Linearly interpolate between all points and from this estimate the CDF of y: $Y=\int_{0}^{x}y(x^{\prime})dx^{\prime}/\int_{0}^{1}y(x^{\prime})dx^{\prime}$.
@@
Once converged (in Figure B this is after $5$ iterations), we construct a grid of $I$ ordered $x$ values from the final $Y(x)$ and convert to the desired set of lags $\Delta t^{i}$} with $\Delta t^{i}=\Delta t_{\mathrm{Min}}+x_{i}^{J-1}(\Delta t_{\mathrm{Max}}-\Delta t_{\mathrm{Min}})$.
@@
Appendix C Optimisation, Preconditioning & Convergence Tests
@@
$\mu_{\mathrm{cont/resp}}$ $=\left(\sum_{i}E_{i}^{-2}\right)^{-1}\sum_{i}{E_{i}^{-2}y_{i}},$ (22)
$\sigma_{\mathrm{cont/resp}}$ $=\left(\sum_{i}E_{i}^{-2}\right)^{-1}\sum_{i}{E_{i}^{-2}(y_{i}-\mu)^{2}}.$
To find a good starting value for $\tau$, we use the ICCF method outlined in Section 3.2 to estimate the auto-covariance function of the continuum. This estimate becomes increasingly less accurate at high autocorrelation times $\delta t$, but drawing some window $\lvert\delta t\rvert<a$ it should follow close to the exponential form in Equation 8. Knowing this, we can say that $\ln\left|\mathrm{ACF}(\delta t)\right|\approx-\lvert\delta t\rvert/\tau+2\ln\left|\sigma\right|$. Drawing some window $\lvert\delta t\rvert<a$ and inverting the $\delta t>0$ side, this logarithm becomes roughly linear with slope $\frac{\partial y}{\partial\delta t}$. Taking a standard linear regression of this function, we can then estimate $\tau$. Precision is not too important here, as the model parameter is $\ln\left|\tau\right|$ in GP fitting and so is insensitive to small deviations.
@@
    Exhibit a severe drop in the peak posterior density compared to the previous slice, which would suggest the slice is in one of the furrows shown in Figure 2.
@@
    An approximate form of (2) where we take only the diagonal elements of $H$, i.e. no skewing / rotating of the axes.
For stability purposes, any negative or zero eigenvalues are fixed to unity in approaches (2) and (3).
Convergence & Uncertainty Estimates
$\Delta\ln\left|Z_{i}\right|=-\nabla f^{T}H^{-1}\nabla f.$ (23)
Conveniently, this also has a natural interpretation in terms of the “closeness to the peak”. Noting that the Hessian is the negative inverse of the covariance matrix, this may also be written $\Delta\phi^{T}H\Delta\phi=\Delta\phi^{T}C^{-1}\Delta\phi$, i.e. the square ‘number of standard deviations’ that the estimated solution is from the peak under the Gaussian approximation. This gives both a measure of the numerical uncertainty in $\ln\left|Z_{i}\right|$ and also a natural measure of “closeness” to the optimum.
$E_{\mathrm{int}}\approx\frac{1}{\sqrt{17}}\lvert Z-Z_{1/2}\rvert$ (24)
For the uncertainty in the SVI Quadrature method, we use Equation 25 but with the $\Delta\ln\left|Z_{i}\right|$ of Equation 23 instead estimated from the standard deviation of the ELBO over multiple SVI iterations assuming a $1/\sqrt{N}$ scaling of the uncertainty.
It is worth emphasising that this represents only the numerical uncertainty arising from convergence of the Laplace Quadrature and SVI Quadrature algorithms, not the uncertainty in the evidence that arises from the Gaussian approximation of the posterior density.
Computational Speed & Tuning
A major feature of LITMUS is that it offers complete Bayesian integrals without exorbitant computational cost. As a part of the scale-testing in Section 5, we measure also the full runtime of the different algorithms, inclusive of null hypothesis model fitting for those methods that can do evidence integrals. The examples here are deliberately under-tuned, with default values for most fitting parameters (see Table 3) but with the number of iterations in the SVI Quadrature ELBO calculations at each slice deliberately set high to ensure convergence.
Shown in Figure C-10, we can see that LITMUS’s various fitting methods are of comparable speed to the JAVELIN-like AEIS with default parameters while still being significantly more reliable (see Section 5). The multimodality in these distributions comes from periods in the fitting of mocks with inconvenient posteriors during which proposals / new live points are difficult to find.
It is difficult to fairly compare the run times of different algorithms due to the long list of tuning parameters for each method. With proper tuning, the SVI Quadrature and Laplace Quadrature algorithms can run markedly faster. Note also that this is not a direct comparison with JAVELIN, only with LITMUS’s example implementation of its fitting method, which is still built in jax and numpyro. In this way Figure C-10 is a comparison of the speed of the algorithms and not the codes themselves.
Figure 10: Histogram of the run-times for the five fitting methods in LITMUS over all mocks, using the fitting parameters described in Table 3. The ICCF method, which requires no matrix inversion owing to its absence of GP fitting, is consistently the fastest. The Laplace Quadrature can perform very fast except for cases where it get stuck optimising at a new test lag when the local optimum changes quickly over the lag axis. The SVI Quadrature has a similar issue, but runs overall somewhat slower.
Laplace Quadrature Parameters
Number of Test lags 128
Grid Bunching Parameter 0.5
Slice Optimisation Tolerance 0.01
Slice Optimisation BFGS Step Size 0.001
Max Optimisation Steps 1000
Slice Discard Log-Density Threshold 10000
SVI Quadrature Parameters
Grid Bunching Parameter 0.5
SVI ELBO Optimisation Steps 1000
SVI Slice Discard ELBO Threshold 1000
ELBO adam Step Size 0.005
Number of ELBO Particles 128
Nest. Samp. Parameters
Number of live points 1000
Max samples 10000
Target evidence uncertainty 0.001
AEIS Parameters
Number of Walkers in Ensemble 256
Total Number of Samples 200000
Warmup Samples 5000
ICCF Parameters
Number of Lags in Grid 512
Number of Bootstraps 512
Number of Times for Light Curve Interpolation 1024
Table 3: Tuning parameters for the fitting methods used in Section 5.